OpenStudy (anonymous):

Which number would you add to one and nine-tenths to give an approximate sum of 6?

OpenStudy (anonymous):

one and 9 tenths is 1.9, so: 1.9 + x = 6 x = 6 - 1.9 x = 4.1 It is unclear what the 'approximate' refers to, but I suspect the asker wants an integer, so say '4'